# Revisting the Impact of Upstream Mergers with Downstream Complements and Substitutes

### v2

@@ 2.4 Implications
Another option would be to use consumer surveys similar to those sometimes used by antitrust authorities to calculate diversion ratios in horizontal merger analyses (Reynolds and Walters, 2008). For instance, in the Somerfield/Morrisons case—a merger between two supermarkets—the UK Competition Commission (now the UK Competition and Markets Authority) asked shoppers “what they would have done had the store been closed” (Reynolds and Walters, 2008, p. 419). These types of questions could be easily adapted to estimate loss ratios by asking shoppers, “What would you do if the store stopped carrying product $i$?”
Finally, while consumer loss ratios may provide practical means for assessing spillover complementarities or substitutabilities, it is crucial to recognize their theoretical limitations. As I formally demonstrate in Section 3.3, in more general settings—encompassing additional products, firms’ pricing decisions, or other types of downstream complementarities and substitutabilities—calculating $\text{CL}_{1}+\text{CL}_{2}-\text{CL}_{12}$ serves as a sufficient statistic for evaluating spillover complementarities or substitutabilities between products 1 and 2 only when the following conditions are simultaneously met:
  1. 1.
    Product spillovers have a common source—for instance, when spillovers depend on how the retailer’s portfolio influences overall consumer traffic. [^14] Otherwise, determining whether products 1 and 2 are complements or substitutes in spillovers becomes significantly more complex. It involves estimating how the removal of these products impacts the individual demands of all remaining products and weighting these effects by the retailer’s profit margins for each product.
  2. 2.
    The retailer does not adjust its prices when losing the merging products, or any price adjustments are minimal (I discuss specific settings where this may hold in that section).
@@ 3.1 The Model
As in the simple model of Section 2, I analyze the effects of a merger between suppliers 1 and 2. I assume that the retailer and suppliers negotiate two-part tariff contracts $(w_{i},t_{i})\in\mathbb{R}^{2}$, where $w_{i}$ is the per-unit wholesale price for product $i$, and $t_{i}$ is a lump-sum fee paid by the retailer to supplier $i$. As before, I use Nash-in-Nash as the bargaining protocol, with $\beta\in(0,1)$ denoting the retailer’s bargaining weight in its negotiations with each supplier. Additionally, I continue to assume that if suppliers 1 and 2 form a conglomerate and either of them fails to reach an agreement with the retailer, the retailer loses access to all the conglomerate’s products.
Under these assumptions, a well-known result in the literature is that negotiated wholesale prices equal marginal costs (O’Brien and Shaffer, 2005). This result reflects the fact that equilibrium wholesale prices are set to maximize total industry profits, with the surplus divided through fixed fees $\bm{t}=(t_{1},\dots,t_{n})$. Thus, without loss of generality, I set $w_{i}=c_{i}$ for all $i=1,\dots,n$ and focus on how the merger affects the negotiated fees.
@@
The main results of this section, presented in the next section, are twofold: (i) demonstrating that demand complementarities or substitutabilities do not always transfer upstream to the retailer’s profit function, and (ii) providing sufficient conditions under which these demand relationships do transfer upstream. However, before presenting the main results, two brief remarks about this more general model are in order.
@@ 3.2 Main Results
(7) $\begin{split}&D_{1}(\bm{p})=(1-p_{1})+b[(1-p_{2})+(1-p_{3})]\\ &D_{2}(\bm{p})=(1-p_{2})+b[(1-p_{1})+(1-p_{3})]\qquad\qquad\text{where $|b|<1$.}\\ &D_{3}(\bm{p})=(1-p_{3})+\gamma\sqrt{(1-p_{1})+(1-p_{2})}\end{split}$
It is clear that if $b>0$ and $\gamma>0$, all three products are strict gross complements, while if $b<0$ and $\gamma<0$, all three products are strict gross substitutes. However, as I show in the proof of the proposition, products 1 and 2 can be strict substitutes for the retailer in terms of profits even when $b>0$ and $\gamma>0$. Similarly, products 1 and 2 can be strict complements for the retailer in terms of profits even when $b<0$ and $\gamma<0$.
@@
As noted in the Related Literature section, the second part of Proposition 3.1 contradicts a formal result from O’Brien and Shaffer (2005). In their analysis, they consider a setting with general demands where $n\geq 2$ upstream suppliers sell their products through a single monopoly retailer, with all products being gross substitutes. On p. 584, they assert that because the products are gross substitutes, they will necessarily be substitutes for the retailer. However, as I now demonstrate, there is a flaw in the proof of this result, which appears on p. 593 of their paper.
I begin by replicating their proof using my notation and then proceed to identify the specific error.
Proof in O’Brien and Shaffer (2005).—  Suppose that products are gross substitutes and the retailer’s pre-optimization profit, $R(\bm{q})\equiv\sum_{i}(P_{i}(\bm{q})-c_{i})q_{i}$, is strictly concave in $\bm{q}$. The claim is that products 1 and 2 are strict substitutes for the retailer in terms of profits, i.e., $\Pi^{*}(1,0,\bm{1}_{-12})-\Pi^{*}(0,0,\bm{1}_{-12})>\Pi^{*}(1,1,\bm{1}_{-12})-\Pi^{*}(0,1,\bm{1}_{-12})$.
Define $M(q_{1},q_{2})\equiv\max_{\bm{q}_{-12}}R(\bm{q})$. Note that $M(q_{1},q_{2})$ is strictly concave in $(q_{1},q_{2})$ since $R(\bm{q})$ is strictly concave in $\bm{q}$, and concavity is preserved under maximization. Let $(q_{1}^{*},q_{2}^{*})$, $q_{1}^{\dagger}$, and $q_{2}^{\dagger}$ be the maximizers of $M(q_{1},q_{2})$, $M(q_{1},0)$, and $M(0,q_{2})$, respectively. Note that by construction,
$\Pi^{*}(1,1,\bm{1}_{-12})=M(q_{1}^{*},q_{2}^{*}),\quad$ $\quad\Pi^{*}(0,1,\bm{1}_{-12})=M(0,q_{2}^{\dagger})$
$\Pi^{*}(1,0,\bm{1}_{-12})=M(q_{1}^{\dagger},0),\,\ \quad$ $\quad\Pi^{*}(0,0,\bm{1}_{-12})=M(0,0)$
Hence, showing that products 1 and 2 are strict substitutes for the retailer in terms of profits is equivalent to showing that $M(q_{1}^{\dagger},0)-M(0,0)>M(q_{1}^{*},q_{2}^{*})-M(0,q_{2}^{\dagger})$. The proof then proceeds as follows:
$\begin{split}M(q_{1}^{*},q_{2}^{*})-M(0,0)&>[M(q_{1}^{*},q_{2}^{*})-M(0,q_{2}^{*})]+[M(q_{1}^{*},q_{2}^{*})-M(q_{1}^{*},0)]\\ &\geq[M(q_{1}^{*},q_{2}^{*})-M(0,q_{2}^{\dagger})]+[M(q_{1}^{*},q_{2}^{*})-M(q_{1}^{\dagger},0)]\end{split}$
where the first (strict) inequality follows due to strict concavity of $M(q_{1},q_{2})$, and the second (weak) inequality by the definition of $q_{1}^{\dagger}$ and $q_{2}^{\dagger}$.
Thus, $M(q_{1}^{*},q_{2}^{*})-M(0,0)>[M(q_{1}^{*},q_{2}^{*})-M(0,q_{2}^{\dagger})]+[M(q_{1}^{*},q_{2}^{*})-M(q_{1}^{\dagger},0)]$, or equivalently, $M(q_{1}^{\dagger},0)-M(0,0)>M(q_{1}^{*},q_{2}^{*})-M(0,q_{2}^{\dagger})$. $\square$
Having provided the proof, I pinpoint the mistake. The problem with the proof is that strict concavity is neither necessary nor sufficient to ensure that $M(q_{1}^{*},q_{2}^{*})-M(0,0)>[M(q_{1}^{*},q_{2}^{*})-M(0,q_{2}^{*})]+[M(q_{1}^{*},q_{2}^{*})-M(q_{1}^{*},0)]$. [^18] For a concrete example, consider $M(q_{1},q_{2})=q_{1}+q_{2}+q_{1}q_{2}-q_{1}^{2}-q_{2}^{2}$. This function is strictly concave in $(q_{1},q_{2})$ but is supermodular in $(q_{1},q_{2})$. Hence, for any $q_{1}>0$ and $q_{2}>0$,$M(q_{1},q_{2})-M(0,0)-[M(q_{1},q_{2})-M(0,q_{2})]-[M(q_{1},q_{2})-M(q_{1},0)]=-q_{1}q_{2}<0$. This demonstrates that strict concavity alone is insufficient to establish the required inequality, as it does not guarantee the necessary relationships between the terms. The correct sufficient condition for this inequality to hold is submodularity of $M(q_{1},q_{2})$, which is unrelated to concavity (Vives, 1999, p. 25).
@@
See Appendix A
An important corollary of this proposition is that under linear demands, products are always complements for the retailer in terms of profits if they are gross complements in the downstream market. This follows from the fact that inverse demands will also be linear in this case and, therefore, weakly supermodular in $\bm{q}$.
@@
Since supermodularity is preserved under maximization (Vives, 1999, p. 26), this implies that $M(q_{1},q_{2})\equiv\max_{\bm{q}_{-12}}R(\bm{q})$ is supermodular in $(q_{1},q_{2})$. Defining $(q_{1}^{*},q_{2}^{*})$, $q_{1}^{\dagger}$, and $q_{2}^{\dagger}$ as the maximizers of $M(q_{1},q_{2})$, $M(q_{1},0)$, and $M(0,q_{2})$, respectively, the supermodularity of $M(q_{1},q_{2})$ is sufficient to ensure that (see the proof of Proposition 3.2 for details):
@@
That is, products 1 and 2 are necessarily complements for the retailer in terms of profits when the retailer carries all other products. The proof of Proposition 3.2 extends this argument to any other combination of products the retailer may carry (i.e., for any $\bm{x}_{-12}=(x_{3},\dots,x_{n})$). This demonstrate that, under the stated conditions, $\Pi^{*}(\bm{x})$ is strictly supermodular in $\bm{x}$.
Unfortunately, a similar result does not exist for downstream substitutabilities. Specifically, products are not necessarily substitutes for the retailer in profits, even if they are all strict gross substitutes and inverse demands are weakly submodular in $\bm{q}$. This highlights an intriguing asymmetry between downstream complementarities and downstream substitutabilities in this context.
To understand the source of this asymmetry, observe that if products are strict gross substitutes and inverse demands are weakly submodular in $\bm{q}$, then $R(\bm{q})$ is also strictly submodular in $\bm{q}$:
@@
3.3 Revisiting the Use of Consumer Loss Ratios
I conclude this section by revisiting the usefulness of consumer loss ratios as a measure of profit spillovers. Specifically, I demonstrate that estimating $\text{CL}_{1}+\text{CL}_{2}-\text{CL}_{12}$, where $\text{CL}_{i}$ is the consumer loss ratio of product $i$, and $\text{CL}_{12}$ is the consumer loss ratio of products 1 and 2 combined, serves as a sufficient statistic for evaluating spillover complementarities between products 1 and 2 to the remaining products in the lineup only when the following conditions are simultaneously met:
  1. 1.
    The retailer does not adjust its prices when losing the merging products, or any price adjustments are minimal.
  2. 2.
    Product spillovers have a common source—for instance, when spillovers depend on how the retailer’s portfolio influences overall consumer traffic.
To start, consider the case where $n=3$. Products 1 and 2 are complements for the retailer in terms of profits, i.e., $\Pi^{*}(1,1,1)-\Pi^{*}(0,1,1)>\Pi^{*}(1,0,1)-\Pi^{*}(0,0,1)$, if and only if:
(8) $\left[\pi_{1}^{*}(1,1,1)-\pi_{1}^{*}(1,0,1)\right]+\left[\pi_{2}^{*}(1,1,1)-\pi_{2}^{*}(0,1,1)\right]\\ \geq\left[\pi_{3}^{*}(1,0,1)+\pi_{3}^{*}(0,1,1)-\pi_{3}^{*}(1,1,1)-\pi_{3}^{*}(0,0,1)\right]$
where $\pi_{i}^{*}(\bm{x})$ denotes the profits the retailer obtains from product $i$ in the optimum when it carries portfolio $\bm{x}=(x_{1},x_{2},x_{3})$, i.e., $\pi_{i}^{*}(\bm{x})=x_{i}D_{i}(\bm{p}^{*}(\bm{x}))$. Although condition (8) is very similar to (3), the key difference is that the retailer’s optimal retail prices now vary with the portfolio it carries. Unfortunately, this implies that consumer loss ratios are no longer a sufficient statistic for product spillovers, as spillovers can also arise from price adjustments.
Nevertheless, if the retailer does not adjust its prices when losing the merging products (or if any price adjustments are minimal), then (8) can be approximately expressed as:
$(p_{1}^{*}-c_{1})\left[D_{1}^{*}(1,1,1)-D_{1}^{*}(0,1,1)\right]+(p_{2}^{*}-c_{2})\left[D_{2}^{*}(1,1,1)-D_{2}^{*}(0,1,1)\right]\\ \geq(p_{3}^{*}-c_{3})\underbrace{\left[D_{3}^{*}(1,0,1)+D_{3}^{*}(0,1,1)-D_{3}^{*}(1,1,1)-D_{3}^{*}(0,0,1)\right]}_{=\text{CL}_{12}-\text{CL}_{1}-\text{CL}_{2}}$
Here, $\bm{p}^{*}\equiv(p_{1}^{*},p_{2}^{*},p_{3}^{*})$ denotes the retailer’s equilibrium prices—which, by assumption, remain unaffected by the negotiation’s outcome—and $D_{i}^{*}(\bm{x})$ denotes the equilibrium demand for product $i$ given retail prices $\bm{p}^{*}$ and portfolio $\bm{x}$ [^19] For instance, $D_{i}^{*}(0,1,1)=D_{i}(p_{1}=\infty,p_{2}^{*},p_{3}^{*})$, whereas $D_{i}^{*}(0,0,1)=D_{i}(p_{1}=\infty,p_{2}=\infty,p_{3}^{*})$ In this case, we recover $\text{CL}_{1}+\text{CL}_{2}-\text{CL}_{12}$ as a sufficient statistic for the profit spillovers of products 1 and 2 over product 3, explaining the necessity of condition 1 above.
To understand the necessity of condition 2, suppose that the retailer always charges $\bm{p}^{*}$ downstream, but the retailer carries $n>3$ products. The generalization of (8) to $n$ products is:
(9) $(p_{1}^{*}-c_{1})[D_{1}^{*}(1,1,\bm{1}_{-12})-D_{1}^{*}(1,0,\bm{1}_{-12})]+(p_{2}^{*}-c_{2})[D_{2}^{*}(1,1,\bm{1}_{-12})-D_{2}^{*}(0,1,\bm{1}_{-12})]\\ >\textstyle\sum_{i\neq 1,2}(p_{i}^{*}-c_{i})[D_{i}^{*}(1,0,\bm{1}_{-12})+D_{i}^{*}(0,1,\bm{1}_{-12})-D_{i}^{*}(1,1,\bm{1}_{-12})-D_{i}^{*}(0,0,\bm{1}_{-12})]$
As expression (9) demonstrates, when there are more than three products, determining whether products 1 and 2 are complements or substitutes in spillovers to the remaining items in the retailer’s lineup becomes significantly more complex. It requires estimating how the removal of these products affects the individual demands of all remaining products, followed by weighting these impacts based on the respective profit margins of each product. Needless to say, this is a daunting task.
However, when product spillovers have a common source, i.e., when $D_{i}^{*}(\bm{x})$ can be written as $D_{i}^{*}(\bm{x})=q_{i}(p_{i}^{*})Q^{*}(\bm{x})$—as in the one-stop-shopping model of Section 2—expression (9) simplifies and consumer loss ratios become useful again:
$(p_{1}^{*}-c_{1})[D_{1}^{*}(1,1,\bm{1}_{-12})-D_{1}^{*}(1,0,\bm{1}_{-12})]+(p_{2}^{*}-c_{2})[D_{2}^{*}(1,1,\bm{1}_{-12})-D_{2}^{*}(0,1,\bm{1}_{-12})]\\ >\textstyle\underbrace{[Q^{*}(1,0,\bm{1}_{-12})+Q^{*}(0,1,\bm{1}_{-12})-Q^{*}(1,1,\bm{1}_{-12})-Q^{*}(0,0,\bm{1}_{-12})]}_{=\mathrm{CL}_{12}-\mathrm{CL}_{1}-\mathrm{CL}_{2}}\sum_{i\neq 1,2}(p_{i}^{*}-c_{i})q_{i}(p_{i}^{*})$
As a final note, while the notion that the retailer might not significantly adjust its retail prices when losing the merging products may initially seem far-fetched, there are several circumstances where this could happen. For example:
  1. (i)
    The Merging Suppliers Products Represent only a Small Subset of the Overall Portfolio. In this case, changes to the retailer’s portfolio will have minimal impact on overall pricing decisions. Ho and Lee (2017) provide empirical evidence that this may be the case in the health insurance market in California.
  2. (ii)
    Vertical Restraints such as RPM. Here, the retailer might be contractually obligated to maintain specific retail prices. While the setting analyzed in this paper cannot rationalize the emergence of vertical restraints, the literature has identified several scenarios where such contractual obligations arise in equilibrium. See Motta (2004, chapter 6) and Rey and Vergé (2008) for detailed discussions.
  3. (iii)
    Unobservability of Retail Prices. In retail environments with search frictions, if retail prices are not observable to consumers before visiting the store and products are unrelated in consumption, prices default to the products’ standard single-product monopoly levels due to the well-known Diamond Paradox (1971). That is,
    $p_{i}^{*}=\operatorname*{arg\,max}_{p_{i}}(p_{i}-c_{i})q_{i}(p),\ \text{for all $i=1,\dots,n$}$
    where $q_{i}(p_{i})$ denotes the demand for product $i$ conditional on the consumer visiting the retailer. Hence, in this case, the retailer’s equilibrium prices remain unaffected by the negotiation’s outcome. For a model exhibiting this property, see Rhodes et al. (2021).
@@ 4 Conclusions
From an antitrust perspective, these insights caution against overgeneralizations that downstream complementarities mitigate competitive concerns while downstream substitutabilities exacerbate them. Such misconceptions may lead to misguided merger enforcement decisions, either blocking welfare-enhancing transactions or prematurely approving mergers with anti-competitive motivations. A reevaluation of merger assessment criteria is essential to account for the mechanisms uncovered in this paper.

### v4

@@ 2.4 Implications
An alternative is to use consumer surveys similar to those sometimes employed by antitrust authorities to calculate diversion ratios in horizontal merger analyses (Reynolds and Walters, 2008). For instance, in the Somerfield/Morrisons case—a merger between two supermarkets—the UK Competition Commission (now the UK Competition and Markets Authority) asked shoppers “what they would have done had the store been closed” (Reynolds and Walters, 2008, p. 419). These types of questions could be easily adapted to estimate loss ratios by asking shoppers, “What would you do if the store stopped carrying product $i$?”
Finally, while consumer loss ratios can provide useful evidence about whether products are complements or substitutes in generating spillovers, it is important to recognize their limitations. In more general environments—including those with additional products or endogenous retail prices—two qualifications arise.
First, estimating consumer loss ratios remains feasible only when spillovers have a common source—for instance, when they depend on how the retailer’s portfolio influences overall consumer traffic. [^12] Otherwise, determining whether products 1 and 2 are complements or substitutes in spillovers becomes significantly more complex. It requires estimating how the removal of each product affects the demand for all remaining products and weighting those effects by the retailer’s profit margins (see the Online Appendix for details). Second, even when estimation is feasible, interpreting $\mathrm{CL}_{1}+\mathrm{CL}_{2}-\mathrm{CL}_{12}$ as a test for whether products are complements or substitutes in generating spillovers is only approximate. The issue is that intermediaries’ retail prices may change when one or both products are removed. As a result, the approximation is most reliable when the retailer does not adjust prices in response to the merger, or when any such adjustments are small. In the Online Appendix, I formalize these qualifications and identify conditions under which retail price responses are likely to be limited.
@@ 3.1 The Model
As in the simple model of Section 2, I analyse the effects of a merger between suppliers 1 and 2. I assume that the retailer and suppliers negotiate two-part tariff contracts $(w_{i},t_{i})\in\mathbb{R}^{2}$, where $w_{i}$ is the per-unit wholesale price for product $i$, and $t_{i}$ is a lump-sum fee paid by the retailer to supplier $i$. As before, I use Nash-in-Nash as the bargaining protocol, with $\beta\in(0,1)$ denoting the retailer’s bargaining weight in its negotiations with each supplier. Additionally, I continue to assume that if suppliers 1 and 2 form a conglomerate and either of them fails to reach an agreement with the retailer, the retailer loses access to all the conglomerate’s products.
Under these assumptions, a well-known result in the literature is that negotiated wholesale prices equal marginal costs (O’Brien and Shaffer, 2005). This result reflects the fact that equilibrium wholesale prices are set to maximise total industry profits, with the surplus divided through fixed fees $\bm{t}=(t_{1},\dots,t_{n})$. Thus, without loss of generality, I set $w_{i}=c_{i}$ for all $i=1,\dots,n$ and focus on how the merger affects the negotiated fees.
@@
The main results of this section, presented next, are twofold: (i) demonstrating that demand complementarities or substitutabilities do not always transfer upstream to the retailer’s profit function, and (ii) providing sufficient conditions under which these demand relationships do transfer upstream. However, before presenting the main results, two brief remarks about this more general model are in order.
@@ 3.2 Main Results
    • Be strict complements for the retailer in terms of profits, even when all products are strict gross substitutes.
    • Be strict substitutes for the retailer in terms of profits, even when all products are strict gross complements.
@@
(7) $\begin{split}&D_{1}(\bm{p})=(1-p_{1})+b[(1-p_{2})+(1-p_{3})]\\ &D_{2}(\bm{p})=(1-p_{2})+b[(1-p_{1})+(1-p_{3})]\\ &D_{3}(\bm{p})=(1-p_{3})+\gamma\sqrt{(1-p_{1})+(1-p_{2})}\end{split}$
It is clear that if $b>0$ and $\gamma>0$, all three products are strict gross complements, while if $b<0$ and $\gamma<0$, all three products are strict gross substitutes. However, as I show in the proof of the proposition, products 1 and 2 are strict substitutes for the retailer in terms of profits when $\gamma=1/2$ and $b\to 0^{+}$, and are strict complements when $\gamma=-1/2$ and $b\to 0^{-}$.
@@
As noted in the Related Literature section, the second part of Proposition 3.1 contradicts a formal result from O’Brien and Shaffer (2005). In their analysis, they consider a setting with general demands where $n\geq 2$ upstream suppliers sell their products through a single monopoly retailer, with all products being gross substitutes. On p. 584, they assert that because the products are gross substitutes, they will necessarily be substitutes for the retailer. However, as I show in the Online Appendix, there is a flaw in the proof of this result, which appears on p. 593 of their paper.
@@
See Appendix A.2. ∎
According to this proposition, when third-party products are present, downstream complementarities transfer upstream only if a stronger condition than gross complementarity is satisfied. Specifically, if inverse demands are interpreted as marginal valuations, the products must be gross complements, and the marginal valuations must be supermodular. [^16] If consumers have quasilinear preferences, gross complementarity is a second-order property of the underlying utility function, while supermodularity of inverse demands is a third-order property. Thus, gross complementarity and supermodular inverse demands are related but distinct properties. An important corollary of this proposition is that under linear demands, products are always complements for the retailer in terms of profits if they are gross complements in the downstream market. This follows from the fact that inverse demands are also linear and therefore weakly supermodular.
@@
Since supermodularity is preserved under maximization (Vives, 1999, p. 26), this implies that $M(q_{1},q_{2})\equiv\max_{\bm{q}_{-12}}R(\bm{q})$ is supermodular in $(q_{1},q_{2})$. Defining $(q_{1}^{*},q_{2}^{*})$, $q_{1}^{\dagger}$, and $q_{2}^{\dagger}$ as the maximisers of $M(q_{1},q_{2})$, $M(q_{1},0)$, and $M(0,q_{2})$, respectively, the supermodularity of $M(q_{1},q_{2})$ is sufficient to ensure that (see the proof of Proposition 3.2 for details):
@@
That is, products 1 and 2 are necessarily complements for the retailer in terms of profits when the retailer carries all other products. The proof of Proposition 3.2 extends this argument to any other combination of products the retailer may carry (i.e., for any $\bm{x}_{-12}=(x_{3},\dots,x_{n})$). This demonstrates that, under the stated conditions $\Pi^{*}(\bm{x})$ is strictly supermodular in $\bm{x}$.
Unfortunately, a similar result does not exist for downstream substitutabilities. Specifically, products are not necessarily substitutes for the retailer in profits, even if they are all strict gross substitutes and inverse demands are weakly submodular in quantities. This highlights an intriguing asymmetry between downstream complementarities and downstream substitutabilities in this context.
To understand the source of this asymmetry, observe that if products are strict gross substitutes and inverse demands are weakly submodular, then $R(\bm{q})$ is also strictly submodular in $\bm{q}$:
@@ 4 Conclusions
From an antitrust perspective, these insights caution against the overgeneralization that downstream complementarities always mitigate competitive concerns while substitutabilities invariably exacerbate them. Such misconceptions may lead to misguided merger enforcement decisions, either blocking welfare-enhancing transactions or prematurely approving mergers with anti-competitive motivations. A reevaluation of merger assessment criteria is essential to account for the mechanisms uncovered in this paper.